### v2

@@ 4.2 Limitations
Despite showing great performance in the task of backlit image enhancement, our models have limitation of non-optimal enhancement of the regions where the loss of information occurs, making restoration ambiguous. Our approach has no generative components and cannot regenerate highly-accurate textures when the information is lost. This leads to sub-optimal performance in such tasks as low-light image enhancement, where images are highly under-exposed. (see supplementary material). Moreover, our models are not trained to exactly match pixels of resulting images with ground-truth data, so some sub-optimal lightening of pixels might occur, which results in non-ideal visual quality.
However, existing generative-based models also have limitations. Many of them cannot be trained in unsupervised setting, and require much time and memory to be trained and used. As a consequence, the resolution of images it can work with is limited.
@@ 5 Conclusion
Based on CLIP-LIT method, we have presented two improved CLIP guidance techniques for backlit image enhancement task, named CLIP-LIT-Latent and RAVE. Our methods show similar or superior performance to that of CLIP-LIT in both supervised and unsupervised settings, while having lower time and space requirements. Furthermore, latent vector constructed by RAVE is interpretable and can be used to find biases in the training data, as well as potentially correct them. Moreover, our proposed approach has several advantages over the current SOTA in supervised backlit image enhancement, Diff-Retinex. First, unlike Diff-Retinex, our method can also be applied to the unsupervised setting. Second, it is more lightweight, requiring less compute and space for training and inference, suitable for images with larger resolution.

### v3

@@ 4.2 Limitations
Despite showing great performance in the task of backlit image enhancement, our models have limitation of non-optimal enhancement of the regions where the loss of information occurs, making restoration ambiguous. Our approach has no generative components and cannot regenerate highly-accurate textures when the information is lost. This leads to sub-optimal performance in such tasks as low-light image enhancement, where images are highly under-exposed (see supplementary material). Moreover, our models are not trained to exactly match pixels of resulting images with ground-truth data, so some sub-optimal lightening of pixels might occur, which results in non-ideal visual quality.
However, existing generative-based models also have limitations. Many of them cannot be trained with unpaired data setting, and require much time and memory to be trained and used. They may also hallucinate detail. As a consequence, the resolution of images it can work with is limited.
@@ 5 Conclusion
Based on CLIP-LIT method, we have presented two improved CLIP guidance techniques for backlit image enhancement task, named CLIP-LIT-Latent and RAVE. Our methods show similar or superior performance to that of CLIP-LIT in both cases of training with paired or unpaired data, while having lower time and space requirements. Furthermore, latent vector constructed by RAVE is interpretable and can be used to find biases in the training data, as well as potentially correct them. Moreover, our proposed approach has several advantages over the current SOTA in backlit image enhancement, Diff-Retinex. First, unlike Diff-Retinex, our method can also be trained using unpaired data. Second, it is more lightweight, requiring less compute and space for training and inference, suitable for images with larger resolution.
Acknowledgements
Funding for this research was provided by a Google-DeepMind PhD Studentship, and work utilised Queen Mary’s Andrena HPC facility, supported by QMUL Research-IT. We also sincerely thank Laida Kushnareva for assisting with images and text composition in the paper.